The molar enthalpy of reaction of

C5H12(g) + 8 O2(g) → 5 CO2(g) + 6 H2O(g)

is -3244.8 kJ/mol for C5H12(g). If 100.0g of oxygen gas react, how much heat is released by the reaction?

Do not include a sign in your answer.

1
Expert's answer
2021-03-23T07:25:25-0400

Solution:

The balanced thermochemical equation:

C5H12(g) + 8O2(g) → 5CO2(g) + 6H2O(g), △H = -3244.8 kJ/mol


The equivalences for this thermochemical equation are:

1 mol C5H12 ⇔ 8 mol O2 ⇔ 5 mol CO2 ⇔ 6 mol H2O ⇔ -3244.8 kJ


Convert the amount of oxygen gas (O2) to moles:

(100.0 g O2) × (1 mol O2 / 32 g O2) = 3.125 mol O2


Use the thermochemical equation to determine the energy change:

(3.125 mol O2) × (3244.8 kJ / 8 mol O2) = 1267.5 kJ


Answer: 1267.5 kJ of heat is released by the reaction.